OpenStudy (jjuden):

A right cone has a base diameter that is 6 cm and a slant height that is 27 cm. What is the surface area of the cone?

OpenStudy (jjuden):

\[A.81\pi cm^2 B.84\pi cm^2 C.90\pi cm^2 D. 162\pi cm^2\]

OpenStudy (anonymous):

B. :)

OpenStudy (clobber02):

lol

OpenStudy (clobber02):

@icecreamman

OpenStudy (clobber02):

somthin came in the mail today

OpenStudy (anonymous):

Uh oh i know whats going down....

OpenStudy (clobber02):

DEEZ NUTZ

OpenStudy (jjuden):

why

OpenStudy (clobber02):

lol idk

OpenStudy (clobber02):

b

OpenStudy (clobber02):

the answer is b

OpenStudy (anonymous):

Clobber02

OpenStudy (clobber02):

what?

OpenStudy (anonymous):

Deez Nutz!

OpenStudy (clobber02):

thats what she said

OpenStudy (anonymous):

Got Em'!

OpenStudy (jjuden):

its ha get em

OpenStudy (clobber02):

i thought it was got heem

OpenStudy (anonymous):

It's HA Got Em'!

OpenStudy (clobber02):

cool got it

OpenStudy (anonymous):

r=radius=1/2 diameter s=slanted hight r=3 π*3*27+π*36 (pi * 3 * 27) + (pi * 36) = 367.56634